(x ^ 2 - 8x + 7)/(x ^ 2 - 6x - 7) / ((x ^ 2 + 7x + 6)/(x ^ 2 - x - 42))
Answer by ikleyn(52778)

.

Factor each trinomial mentally


    x^2 -8x + 7 = (x-1)*(x-7)

    x^2 -6x - 7 = (x-7)*(x+1)

    x^2 +7x + 6 = (x+1)*(x+6)

    x^2 -x - 42 = (x-7)*(x+6)


Then write your expression as the product of fractions and make all necessary reductions.


As simple as 2 x 2 = 4.

Solved (i.e. fully explained)